OpenStudy (anonymous):

If f(x)=x-4 and g(x) = 1 over 2x − 2, find f[g(−6)]

OpenStudy (aravindg):

f(g(x) is simply got but putting g(x) in place of x in f(x)

OpenStudy (anonymous):

g(-6)=1/(-12-2)=1/(-14) f(g(-6))=f(-1/14)=(-1/14)-4

OpenStudy (anonymous):

So it would be -3?

OpenStudy (aravindg):